### v1

@@ 1 Introduction
2 Large Language Models Background
@@ 2.2 Large Multimodal Models
2.3 Parameter-Frozen Paradigm and Parameter-Tuning Paradigm
@@ 3.2.1 Coarse-grained Level
The earliest image-text sentiment classification models were feature-based. In [130], the authors used SentiBank to extract 1200 adjective-noun pairs (ANPs) as visual semantic features and employed SentiStrength [129] to compute text sentiment features for handling multimodal tweet sentiment analysis. In [131], the authors presented a cross-media bag-of-words model to represent the text and image of a Weibo tweet as a unified bag-of-words representation. Then some neural network models showed better performance. In [132, 133], the authors used convolutional neural network (CNN) models to get the representation of text and image. In [134], the authors believed that more detailed semantic information in the image is important and constructed HSAN, a hierarchical semantic attentional network based on image caption for coarse-level multimodal sentiment analysis. MultiSentiNet [135] focused on the correlation between images and text, aggregating the representation of informative words with visual semantic features, objects, and scenes. Considering the mutual influence between image and text, Co-Mem [136] is designed to iteratively model the interactions between visual contents and textual words for multimodal sentiment analysis. In [31], the authors found that images play a supporting role to text in many sentiment detection cases, and proposed VistaNet which instead of using visual information as features only rely on visual information as alignment for pointing out the important sentences of a document using attention. CLMLF [138] applies contrastive learning and data augmentation to align and fuse the token-level features of text and image. In addition to focusing on sentiment, emotions are equally important. In [127], the authors built an image-text emotion dataset, named TumEmo, and further proposed MVAN for multi-modal emotion analysis. In [137], the authors observed that multimodal emotion expressions have specific global features and introduced a graph neural network, proposing an emotion-aware multichannel graph neural network method called MGNNS. MULSER [139] is also a graph-based fusion method that not only investigates the semantic relationship among objects and words respectively, but also explores the semantic relationship between regional objects and global concepts, which has also yielded effective results.
@@
Traditional multimodal sentiment analysis methods often rely on superficial information, lacking the depth provided by contextual world knowledge. This limits their ability to accurately interpret the sentiment conveyed in multimodal content (images and text). WisdoM [141] leverage the contextual world knowledge induced from the LMMs for enhanced multimodal sentiment analysis. The process involves three stages: 1) Prompt Templates Generation: Using ChatGPT to create templates that help LMMs understand the context better. 2) Context Generation: Feeding these templates into LMMs along with the sentence and image to generate rich contextual information. 3) Contextual Fusion: Combining this contextual information with the original sentiment predictions to enhance accuracy, particularly for difficult samples. A training-free module called Contextual Fusion is introduced to minimize noise in the contextual data, ensuring that only relevant information is considered during sentiment analysis. WisdoM significantly outperforms existing state-of-the-art methods in MSED dataset, demonstrating its effectiveness in integrating contextual knowledge for improved sentiment classification. In addition, inspired by the success of textual prompt-based fine-tuning approaches in few-shot scenario, the authors [168] introduce a multi-modal prompt-based fine-tuning approach UP-MPF and the authors [169] propose a prompt-based vision-aware language modeling (PVLM) for multimodal sentiment analysis.
@@
Figure 2: Image-text fine-grained sentiment analysis tasks..
@@ 3.2.2 Fine-grained Level
Inspired by text-based aspect term extraction methods [142, 143, 144], MATE approaches usually view this task as a sequence labeling problem. How to utilize visual information to improve the accuracy of aspect term recognition is the key to this task. Some studies [149, 145] focused on named entity recognition suggest using ResNet encoding to leverage whole image information to enhance the representation of each word. Various neural network-based methods have been developed, including those using recurrent neural networks [145, 146], Transformers [96, 147, 148], and graph neural networks [150]. However, these methods are relatively independent and often focus more on entity information while neglecting the emotional information of the target. Therefore, as research progresses, more scholars in multimodal scenarios are not only extracting aspect terms but also jointly performing corresponding sentiment classification. For example, inspired by the fuzzy span universal information extraction (FSUIE) framework [152] which focuses on the limited span length features in information extraction and proposes fuzzy span loss and fuzzy span attention, DQPSA [151] addresses MATE and MASC under a unified framework as span recognition, dispensing with the complex sequence generation structure.
@@
Different from text-based aspect sentiment classification [153, 154, 155], it is challenging to effectively discover visual sentiment information and fuse it with textual sentiment information. In [77], the authors constructed the Multi-ZOL dataset for the MASC task. This dataset collects and organizes comments about smartphones from the ZOL.com business portal website. At the same time, they proposed a multimodal interactive memory network(MIMN) based on an attention mechanism to capture the information interaction between different modalities. In addition, other researchers [78, 79] have proposed models like the LSTM-based ESAFN model and the Transformer-based TomBERT model for the MASC task, enhancing the interaction of inter-modal and intra-modal sentiment information is the core of these models. Compared with other multimodal tasks such as image and text retrieval, the sentiment annotation used in the MASC task lack strong supervision signals for cross-modal alignment. This issue makes it difficult for MASC models to learn cross-modal interactions and causes models to learn the bias brought by the image. In [32], the authors propose a new method to utilize visual modalities, the image caption generation module in their model undertakes the task of cross-modal alignment. They convert images into text descriptions based on the idea of cross-modal translation. In [98], the authors continued with the idea of modal transformation and employed facial emotions as a supervised signal for learning visual emotions.
As large language models evolve, LLMs and LMMs have adapted to various tasks [156, 157, 158, 159], but their use in the MASC task is still in its initial stages. In [160], the authors attempt to apply the instruction tuning paradigm to the MASC task and utilize the capabilities of LMMs to mitigate the limitations of text and image modality fusion. To address the potential irrelevance between aspects and images, a plug-and-play selector is proposed to autonomously choose the most suitable instruction from the instruction pool, thereby reducing the impact of irrelevant image noise on the final sentiment classification outcome.
@@
As a pioneer, in [161], the authors proposed joint multimodal aspect-sentiment analysis, which jointly performs multimodal aspect term extraction and multimodal aspect sentiment classification. Benefiting from the advancements in visual-language pre-train models, in [80], the authors have designed multimodal sentiment pre-training tasks and develop a unified multimodal encoder-decoder architecture pre-training model for JMASA. In [162], the authors utilize a Cross-Modal Multi-Task Transformer (CMMT) to derive sentiment-aware features for each modality and dynamically control the impact of visual information on textual content during inter-modal interaction. However, the innate semantic gap between visual and language modalities remains a huge challenge for the use of these methods, in [163], the authors believe that the aesthetic attributes of images potentially convey a more profound emotional expression than basic image features and propesd Atlantis. Some scholars [164] have also noticed the impact of image-text pair quality, finding that many studies have overestimated the importance of images due to the presence of many noise images unrelated to text in the datasets. Drawing from the concept of curriculum learning, they proposed a Multi-grained Multi-curriculum Denoising Framework (M2DF), which achieves denoising by adjusting the order of the training data. AOM [165] is designed with an aspect-aware attention module that simultaneously selects text tokens and image blocks semantically related to the aspect to detect semantic and emotional information related to the aspect, thereby reducing noise introduced during the cross-modal alignment process. RNG [166] to simultaneously reduce multi-level modality noise and multigrained semantic gap, design three constraints: (i) Global Relevance Constraint (GR-Con) based on text-image similarity for instance-level noise reduction, (ii) Information Bottleneck Constraint (IB-Con) based on the Information Bottleneck (IB) principle for feature-level noise reduction, and (iii) Semantic Consistency Constraint (SC-Con) based on mutual information maximization in a contrastive learning way for multi-grained semantic gap reduction. To bridge the semantic gap between modal spaces and address the interference of irrelevant visual objects at different scales, in [167], the authors proposed a Multi-level Text-Visual Alignment and Fusion Network (MTVAF).
@@
With the help of LLMs, the JMASA task has also seen further development in recent years. In [170], the authors found that compared to converting MASC into a masked language modeling (MLM) task based on limited sentiment categories with PVLM [169] and UP-MPF [168], MLM is not suitable for JMASA and MATE tasks. They proposed a novel Generative Multimodal Prompt (GMP) model. In [171], the authors explored the potential of using the representative large language model ChatGPT for In-Context Learning (ICL) on the JMASA task. They developed a versatile ICL framework, incorporating zero-shot learning task instructions and expanded it to few-shot learning by adding some demonstration samples to the prompts. Additionally, to enhance the ICL framework’s performance in few-shot learning scenarios, they further developed an entity-aware contrastive learning model to effectively retrieve demonstration samples similar to each test sample.
@@ 3.3.1 Cross-modal Sentiment Semantic Alignment
Attention-based alignment. The attention mechanism has been proven to be an effective method for cross-modal semantic alignment in the field of multimodal learning [179]. Not only can the attention mechanism learn to adapt the alignment relationships for specific tasks through the optimization of task-specific objective functions, but it can also provide a degree of interpretability by outputting attention weights. For example, in the field of image captioning, the attention mechanism focuses on relevant areas when generating text words [180], demonstrating the alignment relationship between words and image regions. Inspired by related research in the multimodal learning field, in [43], the authors proposed using a cross-modal attention mechanism to learn the alignment relationships between pairs of modalities and developed a transformer-based multimodal sentiment analysis model named MulT. The core of the MulT model lies in modeling cross-modal alignment relationships by inserting cross-modal attention layers into the transformer module, allowing dynamic alignment and fusion of fine-grained sentiment information from various modalities. Building on the cross-modal attention mechanism designed in MulT, in [181], the authors introduced the cubic attention mechanism, which generates a three-dimensional attention tensor through parameter computations, representing the alignment information among the three modal representations.
@@
Contrastive learning-based alignment. Contrastive learning achieves cross-modal representation alignment by bringing the representations of positive examples closer together and pushing the representations of negative examples farther apart. A classic model in the field of multimodal learning, CLIP [116], uses contrastive learning to align the semantic representations of text and image modalities, significantly enhancing the quality of image representations and achieving excellent results in tasks such as zero-shot image classification. Inspired by this, the field of audio-image-text sentiment analysis has adopted contrastive learning methods for sentiment semantic representation alignment. In [182], the authors proposed achieving cross-modal emotional semantic alignment by bringing closer the representations of different modalities within the same sample. In [183], the authors suggested using the text-audio and text-image modal information of input samples to predict the corresponding image and audio representations of the samples, then aligning the predicted representations with the actual ones and distancing representations from different samples, thereby aligning the semantic representations of different modalities within the same sample.
@@
Cross-domain transfer learning-based alignment. The field of cross-domain transfer learning primarily studies how to align the sample spaces of target domains with those of source domains so that classifiers trained in the source domains can be directly reused in the target domains. The objectives of this field align broadly with those of cross-modal sentiment representation alignment, hence some studies have explored using cross-domain transfer learning methods for sentiment semantic representation alignment. In [184], considering the rich information content of textual representations, the authors proposed using Deep Canonical Correlation Analysis (DCCA) to align audio and visual representations with textual representations, thereby enhancing the audio and visual representations. In [185], the authors explored using a metric-based domain transfer method, utilizing Central Moment Discrepancy (CMD) to design a loss function that aligns the representations of the three modalities within the same sample. In [186] the authors have employed adversarial learning methods to align sentiment semantic representations across different modalities.

### v2

@@ 1 Introduction
2 Background on Large Language Models
@@ 2.2 Large Multimodal Models
2.3 Usage of Large Language Models
@@ 3.2.1 Coarse-grained Level
The earliest image-text sentiment classification models were feature-based. In [130], the authors used SentiBank to extract 1200 adjective-noun pairs (ANPs) as visual semantic features and employed SentiStrength [129] to compute text sentiment features for handling multimodal tweet sentiment analysis. In [131], the authors presented a cross-media bag-of-words model to represent the text and image of a Weibo tweet as a unified bag-of-words representation. Then some neural network models showed better performance. In [132, 133], the authors used convolutional neural network (CNN) models to get the representation of text and image. In [134], the authors believed that more detailed semantic information in the image is important and constructed HSAN, a hierarchical semantic attentional network based on image caption for coarse-level multimodal sentiment analysis. MultiSentiNet [135] focused on the correlation between images and text, aggregating the representation of informative words with visual semantic features, objects, and scenes. Considering the mutual influence between image and text, Co-Mem [136] is designed to iteratively model the interactions between visual contents and textual words for multimodal sentiment analysis.
In [31], the authors found that images play a supporting role to text in many sentiment detection cases, and proposed VistaNet which instead of using visual information as features only rely on visual information as alignment for pointing out the important sentences of a document using attention. With respect to each image representation $f^{v}_{j}$ , the goal is to learn the attention weights $\beta_{j,i}$ for text representations $f^{t}_{j}$:
$p_{j}=\tanh(W_{p}f^{v}_{j}+b_{p}),$ (5)
$q_{i}=\tanh(W_{q}f^{t}_{i}+b_{q}),$ (6)
$v_{j,i}=V^{T}(p_{j}\odot q_{i}+q_{i}),$ (7)
$\beta_{j,i}=\frac{\exp(v_{j,i})}{\sum_{i}\exp(v_{j,i})}$ (8)
where $W_{p},W_{q},b_{p},b_{q}$ are learnable parameters, firstly, projecting both image representation and text representation onto an attention space followed by a non-linear activation function $\tanh$. Then, let the image projection $p_{j}$ interact with the sentence projection $q_{i}$ in two ways: element-wise multiplication and summation.
CLMLF [138] applies contrastive learning and data augmentation to align and fuse the token-level features of text and image. In addition to focusing on sentiment, emotions are equally important. In [127], the authors built an image-text emotion dataset, named TumEmo, and further proposed MVAN for multi-modal emotion analysis. In [137], the authors observed that multimodal emotion expressions have specific global features and introduced a graph neural network, proposing an emotion-aware multichannel graph neural network method called MGNNS. MULSER [139] is also a graph-based fusion method that not only investigates the semantic relationship among objects and words respectively, but also explores the semantic relationship between regional objects and global concepts, which has also yielded effective results.
Figure 2: Overview of WisdoM framework [171] architecture for coarse-grained image-text sentimen.
@@
Traditional multimodal sentiment analysis methods often rely on superficial information, lacking the depth provided by contextual world knowledge. This limits their ability to accurately interpret the sentiment conveyed in multimodal content (images and text). WisdoM [141] leverage the contextual world knowledge induced from the LMMs for enhanced multimodal sentiment analysis. As shown in Figure 2, the process involves three stages: 1) Prompt Templates Generation: Using ChatGPT to create templates that help LMMs understand the context better. 2) Context Generation: Feeding these templates into LMMs along with the sentence and image to generate rich contextual information. 3) Contextual Fusion: Combining this contextual information with the original sentiment predictions to enhance accuracy, particularly for difficult samples. A training-free module called Contextual Fusion is introduced to minimize noise in the contextual data, ensuring that only relevant information is considered during sentiment analysis. WisdoM significantly outperforms existing state-of-the-art methods in MSED dataset, demonstrating its effectiveness in integrating contextual knowledge for improved sentiment classification. In addition, inspired by the success of textual prompt-based fine-tuning approaches in few-shot scenario, the authors [168] introduce a multi-modal prompt-based fine-tuning approach UP-MPF and the authors [169] propose a prompt-based vision-aware language modeling (PVLM) for multimodal sentiment analysis.
@@
Figure 3: Image-text fine-grained sentiment analysis tasks.
@@ 3.2.2 Fine-grained Level
Inspired by text-based aspect term extraction methods [142, 143, 144], MATE approaches usually view this task as a sequence labeling problem. How to utilize visual information to improve the accuracy of aspect term recognition is the key to this task. Some studies [149, 145] focused on named entity recognition suggest using ResNet encoding to leverage whole image information to enhance the representation of each word. Various neural network-based methods have been developed, including those using recurrent neural networks [145, 146], Transformers [96, 147, 148], and graph neural networks [150]. Conditional Random Fields (CRF) are widely used in sequence labeling tasks because CRF considers the correlations between labels in neighborhoods. For example, an adjective has a greater probability of being followed by a noun than a verb in POS tagging task. Using $Y=(y_{1},y_{2},…,y_{n})$ represents a generic sequence of labels for input $S$. Given sequence $S$, all the possible label sequences $Y$ can be calculated by the following equation:
$p(Y|S)=\frac{\prod_{i=1}^{n}\Omega(y_{i-1},y_{i},X)}{\sum_{y^{{}^{\prime}}\in Y}\prod_{i=1}^{n}\Omega(y^{{}^{\prime}}_{i-1},y^{{}^{\prime}}_{i},X)},$ (9)
where $\Omega(y_{i-1},y_{i},X)$ and $\Omega(y^{{}^{\prime}}_{i-1},y^{{}^{\prime}}_{i},X)$ are potential functions.
However, these methods are relatively independent and often focus more on entity information while neglecting the emotional information of the target. Therefore, as research progresses, more scholars in multimodal scenarios are not only extracting aspect terms but also jointly performing corresponding sentiment classification.
Figure 4: Illustration of exact boundary and fuzzy boundary.
Figure 5: Overview of DQPSA model [151] architecture.
For example, inspired by the fuzzy span universal information extraction (FSUIE) framework [152] which focuses on the limited span length features in information extraction and proposes fuzzy span loss and fuzzy span attention, DQPSA [151] addresses MATE and MASC under a unified framework as span recognition, dispensing with the complex sequence generation structure. As shown in Figure 5, To capture the semantical relevance of start and end boundaries within a span, DQPSA introduce the idea of Energy Based Model [216] to give better span scores and propose the Energy based Pairwise Expert that predicts span based on pairwise stability.
@@
Different from text-based aspect sentiment classification [153, 154, 155], it is challenging to effectively discover visual sentiment information and fuse it with textual sentiment information. In [77], the authors constructed the Multi-ZOL dataset for the MASC task. This dataset collects and organizes comments about smartphones from the ZOL.com business portal website. At the same time, they proposed a multimodal interactive memory network(MIMN) based on an attention mechanism to capture the information interaction between different modalities. In addition, other researchers [78, 79] have proposed models like the LSTM-based ESAFN model and the Transformer-based TomBERT model for the MASC task, enhancing the interaction of inter-modal and intra-modal sentiment information is the core of these models. The TomBERT model treat the hidden states of the target aspect $A$ as queries, and the regional image features $H^{V}$ as keys and values, such that the target is leveraged to guide the model to align it with the appropriate regions.
$ATT_{i}(A,H^{V})=softmax(\frac{[W_{Q_{i}}A]^{T}[W_{K_{i}}H^{V}]}{\sqrt{d//m}})[W_{V_{i}}H^{V}]^{T},$ (10)
where ${W_{Q_{i}},W_{K_{i}},W_{V_{i}}}\in R^{d/m\times d}$ are parameters, $m$ represent the attention head number, $d$ is the input embeddings dimension.
Compared with other multimodal tasks such as image and text retrieval, the sentiment annotation used in the MASC task lack strong supervision signals for cross-modal alignment. This issue makes it difficult for MASC models to learn cross-modal interactions and causes models to learn the bias brought by the image. In [32], the authors propose a new method to utilize visual modalities, the image caption generation module in their model undertakes the task of cross-modal alignment. They convert images into text descriptions based on the idea of cross-modal translation.
$C=Caption\_Transformer(V),$ (11)
where $C$ and $Caption_{t}ransformer$ denote the output image captions and the transformer-based image caption generator, respectively. In sentence-pair classification mode, input to pre-trained language model takes the sentence-pair form
$[CLS]w_{1}^{T},w_{2}^{T}...w_{T.len}^{T}[SEP]w_{1}^{C},w_{2}^{C}...w_{C.len}^{C}[PAD]...[PAD],$ (12)
where $w_{i}^{T}$ are the tokens of the input text, and $w_{i}^{C}$ are the tokens of the image caption. In [98], the authors continued with the idea of modal transformation and employed facial emotions as a supervised signal for learning visual emotions.
Figure 6: Overview of A2II model [160] architecture for MASC task.
As large language models evolve, LLMs and LMMs have adapted to various tasks [156, 157, 158, 159], but their use in the MASC task is still in its initial stages. In [160], the authors attempt to apply the instruction tuning paradigm to the MASC task and utilize the capabilities of LMMs to mitigate the limitations of text and image modality fusion. As shown in Figure 6, to address the potential irrelevance between aspects and images, a plug-and-play selector is proposed to autonomously choose the most suitable instruction from the instruction pool, thereby reducing the impact of irrelevant image noise on the final sentiment classification outcome.
@@
As a pioneer, in [161], the authors proposed joint multimodal aspect-sentiment analysis, which jointly performs multimodal aspect term extraction and multimodal aspect sentiment classification. Since it is a joint task with aspect terms extraction and aspect sentiment classification, the authors calculate two different sets of loss simultaneously as follows:
$L=-\sum^{k}_{i=1}y_{i}^{s}\log p_{i}^{str}-\sum^{k}_{i=1}y_{i}^{e}\log p_{i}^{end}-\sum^{m}_{t=1}\sum^{\epsilon}_{i=1}y_{ti}^{p}\log p_{ti}^{p}$ (13)
where $y^{s},y^{e},y^{p}$ are one-hot labels indicating golden start, end positions, true sentiment polarity separately, and $a,m$ are the number of sentence tokens, aspects respectively.
@@
Benefiting from the advancements in visual-language pre-train models, in [80], the authors have designed multimodal sentiment pre-training tasks and develop a unified multimodal encoder-decoder architecture pre-training model for JMASA. In [162], the authors utilize a Cross-Modal Multi-Task Transformer (CMMT) to derive sentiment-aware features for each modality and dynamically control the impact of visual information on textual content during inter-modal interaction. However, the innate semantic gap between visual and language modalities remains a huge challenge for the use of these methods, in [163], the authors believe that the aesthetic attributes of images potentially convey a more profound emotional expression than basic image features and propesd Atlantis. Some scholars [164] have also noticed the impact of image-text pair quality, finding that many studies have overestimated the importance of images due to the presence of many noise images unrelated to text in the datasets. Drawing from the concept of curriculum learning, they proposed a Multi-grained Multi-curriculum Denoising Framework (M2DF), which achieves denoising by adjusting the order of the training data. AOM [165] is designed with an aspect-aware attention module that simultaneously selects text tokens and image blocks semantically related to the aspect to detect semantic and emotional information related to the aspect, thereby reducing noise introduced during the cross-modal alignment process. RNG [166] to simultaneously reduce multi-level modality noise and multigrained semantic gap, design three constraints: (i) Global Relevance Constraint (GR-Con) based on text-image similarity for instance-level noise reduction, (ii) Information Bottleneck Constraint (IB-Con) based on the Information Bottleneck (IB) principle for feature-level noise reduction, and (iii) Semantic Consistency Constraint (SC-Con) based on mutual information maximization in a contrastive learning way for multi-grained semantic gap reduction. To bridge the semantic gap between modal spaces and address the interference of irrelevant visual objects at different scales, in [167], the authors proposed a Multi-level Text-Visual Alignment and Fusion Network (MTVAF).
Figure 7: Overview of versatile ICL framework [171] architecture for JMASA task.
@@
With the help of LLMs, the JMASA task has also seen further development in recent years. In [170], the authors found that compared to converting MASC into a masked language modeling (MLM) task based on limited sentiment categories with PVLM [169] and UP-MPF [168], MLM is not suitable for JMASA and MATE tasks. They proposed a novel Generative Multimodal Prompt (GMP) model. In [171], the authors explored the potential of using the representative large language model ChatGPT for In-Context Learning (ICL) on the JMASA task. As shown in Figure 7, they developed a versatile ICL framework, incorporating zero-shot learning task instructions and expanded it to few-shot learning by adding some demonstration samples to the prompts. Additionally, to enhance the ICL framework’s performance in few-shot learning scenarios, they further developed an entity-aware contrastive learning model to effectively retrieve demonstration samples similar to each test sample.
@@ 3.3.1 Cross-modal Sentiment Semantic Alignment
Attention-based alignment. The attention mechanism has been proven to be an effective method for cross-modal semantic alignment in the field of multimodal learning [179]. Not only can the attention mechanism learn to adapt the alignment relationships for specific tasks through the optimization of task-specific objective functions, but it can also provide a degree of interpretability by outputting attention weights. For example, in the field of image captioning, the attention mechanism focuses on relevant areas when generating text words [180], demonstrating the alignment relationship between words and image regions. Inspired by related research in the multimodal learning field, in [43], the authors proposed using a cross-modal attention mechanism to learn the alignment relationships between pairs of modalities and developed a transformer-based multimodal sentiment analysis model named MulT. The core of the MulT model lies in modeling cross-modal alignment relationships by inserting cross-modal attention layers into the transformer module, allowing dynamic alignment and fusion of fine-grained sentiment information from various modalities. First, use $\alpha$ and $\beta$ to represent two different modalities, define the $\mathrm{Query}$s as $Q_{\alpha}=X_{\alpha}W_{Q_{\alpha}}$, $\mathrm{Key}$s as $K_{\beta}=X_{\beta}W_{K_{\beta}}$, and $\mathrm{Value}$s as $V_{\beta}=X_{\beta}W_{V_{\beta}}$, where $W_{Q_{\alpha}}\in\mathbb{R}^{d_{\alpha}\times d_{k}},W_{K_{\beta}}\in\mathbb{R}^{d_{\beta}\times d_{k}}$ and $W_{V_{\beta}}\in\mathbb{R}^{d_{\beta}\times d_{v}}$ are weights. The latent adaptation from $\beta$ to $\alpha$ is presented as the crossmodal attention $Y_{\alpha}:=\text{CM}_{\beta\rightarrow\alpha}(X_{\alpha},X_{B})\in\mathbb{R}^{T_{\alpha}\times d_{v}}$:
$Y_{\alpha}=\text{CM}_{\beta\rightarrow\alpha}(X_{\alpha},X_{\beta})=\text{softmax}\left(\frac{Q_{\alpha}K_{\beta}^{\top}}{\sqrt{d_{k}}}\right)V_{\beta}=\text{softmax}\left(\frac{X_{\alpha}W_{Q_{\alpha}}W_{K_{\beta}}^{\top}X_{\beta}^{\top}}{\sqrt{d_{k}}}\right)X_{\beta}W_{V_{\beta}}.$ (14)
Note that $Y_{\alpha}$ has the same length as $Q_{\alpha}$, but is meanwhile represented in the feature space of $V_{\beta}$. Specifically, the scaled (by $\sqrt{d_{k}}$) softmax computes a score matrix $\mathrm{softmax\,}(\cdot)\in\mathbb{R}^{T_{\alpha}\times T_{\beta}}$, whose $(i,j)$-th entry measures the attention given by the $i$-th time step of modality $\alpha$ to the $j$-th time step of modality $\beta$. Hence, the $i$-th time step of $Y_{\alpha}$ is a weighted summary of $V_{\beta}$, with the weight determined by $i$-th row in $\mathrm{softmax}(\cdot)$.
Building on the cross-modal attention mechanism designed in MulT, in [181], the authors introduced the cubic attention mechanism, which generates a three-dimensional attention tensor through parameter computations, representing the alignment information among the three modal representations.
@@
Contrastive learning-based alignment. Contrastive learning achieves cross-modal representation alignment by bringing the representations of positive examples closer together and pushing the representations of negative examples farther apart. A classic model in the field of multimodal learning, CLIP [116], uses contrastive learning to align the semantic representations of text and image modalities, significantly enhancing the quality of image representations and achieving excellent results in tasks such as zero-shot image classification. Inspired by this, the field of audio-image-text sentiment analysis has adopted contrastive learning methods for sentiment semantic representation alignment. In [182], the authors proposed achieving cross-modal emotional semantic alignment by bringing closer the representations of different modalities within the same sample. In [183], the authors suggested using the text-audio and text-image modal information of input samples to predict the corresponding image and audio representations of the samples, then aligning the predicted representations with the actual ones and distancing representations from different samples, thereby aligning the semantic representations of different modalities within the same sample. The proposed model contains two key modules, the uni-modal coding drive the model to focus on informative features which then implicitly filter out inherent noise and produces robust and effective uni-modal representation for acoustic and visual modalities.
Given a batch set $F_{uni}=\{F_{u}^{0},F_{u}^{1},...,F_{u}^{n-1}\}$, noted that there is a single positive key $F^{\textit{\textdagger}}_{u}$ (as $k^{+}$) that each encoded query $F_{u}^{i}$ (as $q$, $i\in[1,n]$) matches, while the other representations $F_{u}^{j}$ ($j\in[0,n]$ and $j\neq i$) in the same batch are considered as negative key samples $k^{-}$. With the similarity measured by dot product, the uni-modal instance contrastive loss $\mathcal{L}_{uni}$ in:
$\mathcal{L}^{u}_{uni}\triangleq-\log\frac{\exp(q\cdot k^{+}/\tau)}{\sum_{i=1}^{n}\exp(q\cdot q^{i}/\tau)}=-\mathop{\mathbb{E}}\limits_{F_{uni}}\left[\log\frac{\exp(F_{u}\cdot F^{\textit{\textdagger}}_{u}/\tau)}{\sum_{i=1}^{n}\exp(F_{u}\cdot F_{u}^{i}/\tau)}\right]$ (15)
where $\tau$ is a temperature hyper-parameter that controls the probability distribution over distinct instances. Due to $u\in\{a,v\}$, the final uni-modal instance contrastive loss $\mathcal{L}_{uni}=\mathcal{L}^{a}_{uni}+\mathcal{L}^{v}_{uni}$.
The cross-modal prediction captures commonalities among different modalities and outputs predictive representation full of interaction dynamics. Each query $q$ has a corresponding key as $k^{+}$ while the other representations in the same batch are seen as $k^{-}$. Similar with uni-modal instance contrastive loss $\mathcal{L}_{uni}$, the cross-modal instance contrastive loss $\mathcal{L}_{cross}$ is presented as:
$\mathcal{L}_{cross}\triangleq-\mathop{\mathbb{E}}\limits_{F_{cross}}\left[\log\frac{\exp(F_{c}\cdot F^{+}_{c}/\tau)}{\sum_{i=1}^{n}\exp(F_{c}\cdot F_{c}^{i}/\tau)}\right]$ (16)
where $F_{c}\backslash F_{c}^{+}\in\{P_{u},G_{u}\}$, $P_{u}$ represent the prediction while $G_{u}$ represent the target, $u\in\{a,v\}$ and $F_{cross}=\{F_{c}^{1},...,F_{c}^{n}\}$.
@@
Cross-domain transfer learning-based alignment. The field of cross-domain transfer learning primarily studies how to align the sample spaces of target domains with those of source domains so that classifiers trained in the source domains can be directly reused in the target domains. The objectives of this field align broadly with those of cross-modal sentiment representation alignment, hence some studies have explored using cross-domain transfer learning methods for sentiment semantic representation alignment. In [184], considering the rich information content of textual representations, the authors proposed using Deep Canonical Correlation Analysis (DCCA) to align audio and visual representations with textual representations, thereby enhancing the audio and visual representations. In [185], the authors explored using a metric-based domain transfer method, utilizing Central Moment Discrepancy (CMD) to design a loss function that aligns the representations of the three modalities within the same sample. The overall learning of the model is performed by minimizing:
$\mathcal{L}=\mathcal{L}_{\text{task}}+\alpha\,\mathcal{L}_{\text{sim}}+\beta\,\mathcal{L}_{\text{diff}}+\gamma\,\mathcal{L}_{\text{recon}}$ (17)
Here, $\alpha,\beta,\gamma$ are the interaction weights that determine the contribution of each regularization component to the overall loss $\mathcal{L}$. Minimizing the similarity loss $\mathcal{L}_{\text{sim}}$ reduces the discrepancy between the shared representations of each modality. This helps the common cross-modal features to be aligned together in the shared subspace. Difference Loss $\mathcal{L}_{\text{diff}}$ is to ensure that the modality-invariant and -specific representations capture different aspects of the input. As the difference loss is enforced, there remains a risk of learning trivial representations by the modality-specific encoders. To avoid this situation, the authors add a reconstruction loss $\mathcal{L}_{\text{recon}}$ that ensures the hidden representations to capture details of their respective modality. The task-specific loss $\mathcal{L}_{\text{task}}$ estimates the quality of prediction during training.
In [186] the authors have employed adversarial learning methods to align sentiment semantic representations across different modalities.